f n=450 and p' (p-prime) = 0.06, construct a 99% confidence interval.

Give your answers to three decimals.

< p <
Answer #1

Solution :

Given that,

n = 450

\hat p =0.06

1 -\hat p = 1 - 0.06 = 0.94

At 99% confidence level the z is ,

\alpha  = 1 - 99% = 1 - 0.99 = 0.01

\alpha / 2 = 0.01 / 2 = 0.005

Z\alpha/2 = Z0.005 = 2.576

Margin of error = E = Z\alpha / 2 * \sqrt((\hat p * (1 - \hat p )) / n)

= 2.576 * (\sqrt((0.06 * 0.94) / 450) = 0.029

A 99 % confidence interval for population proportion p is ,

\hat p - E < P <\hat p + E

0.06 - 0.029 < p < 0.06 + 0.029

0.031< p < 0.089

The 99% confidence interval for the population proportion p is : ( 0.031< p < 0.089)